Determine the square root of 1296
1
Expert's answer
2020-02-10T09:47:45-0500

"\sqrt{1296} = 36"

We may prove it if we know simplier square root values.

For example, we know that 22 = 4 and 1296 may be divided by 4(because its last 2 digits (96) may be divided by 4).

So, "\sqrt{1296} = \sqrt{\frac{1296}{4}} * \sqrt{4} = 2 * \sqrt{324}"

324 may also be divided by 4 (as its last 2 digits (24) may be divided by 4).

Thus, "2 * \sqrt{324} = 2* \sqrt{\frac{324}{4}} * \sqrt{4} = 4 * \sqrt{81}"

As we know, "\sqrt{81} = 9" , so "\sqrt{1296} = 4 * 9 = 36" .